OpenStudy (anonymous):

find a trigonometric representation of the solutions in C (complex numbers) of the equation x^3 -1 = 0

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

x^3=1 x^3=cos0+isin0 x=(cos (2k pi +0)+ isin(2k pi+0))^(1/3) = cos((2k pi)/3) + i sin ((2k pi)/3) (DeMoivre's thm) where k=0,1,2.....